# Core-Set Selection for Data-efficient Land Cover Segmentation

### v1

@@ IV-B Implementation Details
We pre-process the aforementioned datasets by tiling them into non-overlapping $256\times 256$ patches that are used in all subsequent steps.
@@
All proposed methods [^1] The code has been made publicly available at https://github.com/keillernogueira/data-centric-rs-classification/ were implemented using Pytorch. During training, we use the following hyper-parameters: 100 training epochs, AdamW [28] as optimizer, learning rate of 0.001, and batch size of 64 for the DFC2022 dataset and 32 for the other datasets.
@@
IV-C Experimental Protocol
For the DFC2022 dataset, 90% of the data originally released for the data fusion contest was made available to be ranked by the proposed core-set algorithms, while the remaining 10% was used for validation.
For Vaihingen and Potsdam datasets, we followed a protocol similar to the one proposed by [29]. Specifically, for the Vaihingen dataset, 11 images originally released for the contest were made available for the proposed core-set techniques, and 5 images (with IDs 11, 15, 28, 30, 34) were employed for validation. For the Potsdam dataset, 18 images released for the contest were made available for the proposed techniques, whereas 6 images (with IDs 02_12, 03_12, 04_12, 05_12, 06_12, 07_12) were used for validation.
For all datasets, the validation is only employed to assess the training of the U-Net models, after the selection of the core-set. The final evaluation of the trained U-Net models is done using the original test set of each dataset. The overall performance of each method is measured by the mean Intersection over Union (mIoU) across all segmentation classes and averaged over three different model training runs.
@@ V-A Quantitative Results
To compare the performance of the scoring functions, we train and test the U-Net model (using the configuration described in Section IV-B) on the top $1\%$, $5\%$, $10\%$, $25\%$, $50\%$, and $75\%$ ranked patches from each method, and an identically sized random selection of patches for the baseline method. Further, we train the U-Net model on 100% of the data as a strong baseline. To account for potential variability due to randomness, three models are trained for each approach and subset size, using the same selected samples (per subset) and hyperparameters. We use a paired t-test with $\alpha=0.05$ to evaluate statistically significant differences in results across methods.
Table I shows the results for each method across the three datasets and different core-set sizes. Overall, the proposed approaches outperform the baseline for all datasets, even when the latter is trained using 100% of the available data. In most cases, the proposed approaches outperform the baseline even when using substantially fewer training samples. For instance, in the Vaihingen dataset, the FA/CB Hybrid at 10% achieved a higher mIoU score than the baseline on the top-ranked 25% of the data. This shows the ability of the proposed techniques to select representative samples (core-set) for training.
@@
Finally, we also observe that each dataset has a different IoU convergence rate. DFC2022 has the most label noise (as evidenced by performance degradation of models in later splits), and the best-performing methods reached close to peak performance by utilizing only 10% of the data, noticeably outperforming training on 100% of the data. This rapid convergence suggests that for datasets with specific characteristics (e.g., redundancy, noise), a relatively small core-set can be as effective, or even more effective, than the full dataset. In contrast, on datasets with less label noise, such as Vaihingen and Potsdam, the performance continues to gradually improve as the number of training samples increases. However, even in this case, the proposed methods are able to outperform the baseline trained with 100% of the samples, demonstrating the importance of selecting a core set to deal with relevant issues such as noise, representativeness, and so on. In all cases, at least one proposed approach outperformed the baseline trained using 100% of the available samples, thus showcasing the advantage of the core-set selection for remote sensing semantic segmentation tasks, reducing training time while improving performance, as well as the ability of the proposed approaches to select such sets.
TABLE I: Results (mIoU) across different sized subsets of data (1%, 5%, 10%, 25%, 50%, 75%, and 100%) for DFC2022, Vaihingen, and Potsdam datasets. Underlined values indicate the results that outperformed the corresponding baseline per training percentage (statistically significant paired t-test at $\alpha=0.05$). Bold values represent the best results overall for the dataset.
@@
1% 5% 10% 25% 50% 75% 100%
Random baseline 10.02 $\pm$ 0.37 11.02 $\pm$ 0.23 11.72 $\pm$ 0.12 12.38 $\pm$ 0.29 12.68 $\pm$ 0.11 12.02 $\pm$ 0.11 12.29 $\pm$ 0.12
Label Complexity (LC) 10.54 $\pm$ 0.63 12.04 $\pm$ 0.20 12.68 $\pm$ 0.20 12.65 $\pm$ 0.03 12.46 $\pm$ 0.16 11.73 $\pm$ 0.76 -
Feature Diversity (FD) 10.71 $\pm$ 0.71 11.57 $\pm$ 0.44 11.77 $\pm$ 0.49 11.96 $\pm$ 0.05 12.17 $\pm$ 0.29 11.86 $\pm$ 0.27 -
LC/FD Hybrid 10.27 $\pm$ 0.26 12.43 $\pm$ 0.25 12.09 $\pm$ 0.50 12.54 $\pm$ 0.14 12.24 $\pm$ 0.34 12.28 $\pm$ 0.35 -
Feature Activation (FA) 9.14 $\pm$ 0.54 10.85 $\pm$ 0.21 10.86 $\pm$ 0.37 11.63 $\pm$ 0.18 12.32 $\pm$ 0.12 12.45 $\pm$ 0.26 -
Class Balance (CB) 9.12 $\pm$ 0.52 10.69 $\pm$ 0.90 11.01 $\pm$ 0.58 10.58 $\pm$ 0.19 11.95 $\pm$ 0.22 11.99 $\pm$ 0.19 -
FA/CB Hybrid 10.13 $\pm$ 0.10 10.14 $\pm$ 0.25 10.82 $\pm$ 0.22 11.98 $\pm$ 0.61 11.62 $\pm$ 0.07 12.10 $\pm$ 0.22 -
@@
1% 5% 10% 25% 50% 75% 100%
Random baseline 36.20 $\pm$ 1.36 42.13 $\pm$ 0.52 50.72 $\pm$ 2.02 50.37 $\pm$ 2.97 55.82 $\pm$ 1.74 59.57 $\pm$ 0.64 59.95 $\pm$ 0.50
Label Complexity (LC) 33.00 $\pm$ 4.03 45.80 $\pm$ 3.62 50.23 $\pm$ 4.33 53.75 $\pm$ 0.89 55.53 $\pm$ 0.41 58.70 $\pm$ 2.06 -
Feature Diversity (FD) 33.67 $\pm$ 3.38 42.15 $\pm$ 0.57 44.87 $\pm$ 0.39 48.92 $\pm$ 3.60 56.07 $\pm$ 1.92 58.63 $\pm$ 0.50 -
LC/FD Hybrid 33.21 $\pm$ 3.50 40.45 $\pm$ 0.64 43.99 $\pm$ 0.42 51.15 $\pm$ 5.23 57.76 $\pm$ 0.46 58.97 $\pm$ 0.88 -
Feature Activation (FA) 35.20 $\pm$ 1.22 43.61 $\pm$ 0.70 51.44 $\pm$ 0.69 53.44 $\pm$ 0.09 56.85 $\pm$ 0.60 60.58 $\pm$ 0.66 -
Class Balance (CB) 40.92 $\pm$ 3.10 44.80 $\pm$ 1.23 51.03 $\pm$ 3.62 55.42 $\pm$ 0.83 58.36 $\pm$ 1.04 60.58 $\pm$ 0.30 -
FA/CB Hybrid 36.88 $\pm$ 1.39 49.06 $\pm$ 1.04 53.56 $\pm$ 0.25 56.92 $\pm$ 0.25 58.96 $\pm$ 0.67 60.43 $\pm$ 0.75 -
@@
1% 5% 10% 25% 50% 75% 100%
Random baseline 60.89 $\pm$ 0.22 66.27 $\pm$ 0.88 68.21 $\pm$ 0.67 72.28 $\pm$ 0.66 74.56 $\pm$ 0.45 78.42 $\pm$ 0.33 78.00 $\pm$ 0.27
Label Complexity (LC) 60.07 $\pm$ 2.33 68.68 $\pm$ 0.39 70.28 $\pm$ 0.45 75.12 $\pm$ 0.52 77.23 $\pm$ 0.79 79.15 $\pm$ 1.23 -
Feature Diversity (FD) 59.78 $\pm$ 0.69 66.56 $\pm$ 0.35 69.37 $\pm$ 0.25 73.77 $\pm$ 0.54 75.42 $\pm$ 1.34 78.35 $\pm$ 1.21 -
LC/FD Hybrid 53.70 $\pm$ 6.05 68.25 $\pm$ 0.16 71.38 $\pm$ 0.34 74.87 $\pm$ 0.54 76.58 $\pm$ 0.56 78.72 $\pm$ 0.37 -
Feature Activation (FA) 56.86 $\pm$ 1.47 65.75 $\pm$ 0.63 68.21 $\pm$ 0.58 72.89 $\pm$ 1.03 74.70 $\pm$ 0.52 79.58 $\pm$ 1.56 -
Class Balance (CB) 49.89 $\pm$ 1.17 58.26 $\pm$ 0.39 61.94 $\pm$ 0.68 70.40 $\pm$ 0.30 75.29 $\pm$ 1.82 79.64 $\pm$ 0.48 -
FA/CB Hybrid 55.29 $\pm$ 1.49 61.72 $\pm$ 0.42 64.62 $\pm$ 0.40 71.06 $\pm$ 0.89 75.42 $\pm$ 0.78 78.49 $\pm$ 0.77 -
In addition to the gains in terms of performance, Table II reports the training time per epoch, in seconds, by a machine with an Intel Xeon E5-2695 v4 “Broadwell” CPU, 128GB of RAM memory, Nvidia P100 ”Pascal” GPU with 16GB of memory, under an 11.2 CUDA version, and Red Hat Enterprise Linux system release 8.2. As all models were trained for 100 epochs, this allows us to contextualize computational savings from the core-set methods. For example, with the DFC2022 dataset, the best-performing result (using 10% of the data) trained for 2 days and 2 hours less than the 100% training baseline while achieving a better test set performance.
@@ V-B Qualitative Results
To better analyze and compare the methods’ outputs, we include visualizations of the average generated rankings, as can be seen in Figure 5. To generate these visualizations, the scores produced by the different methods are first averaged per patch and then ranked, thus making the most frequently selected patches across the methods appear at the top. Additionally, the standard deviation is calculated to capture the variability and provide insight into the approaches’ consistency. Furthermore, to allow for a better analysis, examples of the most and least frequently selected samples across all methods are presented in Figure 6.
@@
Figure 4: Visualizations of the methods’ rankings. The line represents the average rank position for each patch across all proposed approaches, while the shaded area represents the standard deviation.
@@
Figure 5: Visualizations of the methods’ rankings. The line represents the average score for each patch across dall proposed approaches, while the shaded area represents the average standard deviation.
For all datasets, it is possible to observe that the approaches exhibit significant consistency, frequently assigning more importance to a specific (core) set of high-complexity samples that are consistently selected across methods. Furthermore, the methods also tend to agree on the least important samples, assigning lower scores to low-complexity patches, indicating that the explored datasets contain a subset of non-representative or noisy samples that either contribute minimally to the overall performance or, in some cases, may even degrade it. Overall, the ability to select the core set, along with the identification of less valuable samples, highlights the robustness and efficiency of the proposed techniques in distinguishing between high- and low-quality data, thereby resulting in better performance and training time.
@@
Figure 6: Examples of the most frequently selected samples (first three columns) and the least frequently selected instances (last three columns) considering the average ranking of all methods.
@@ VI Conclusions
In this paper, we propose several core-set selection approaches for remote sensing image segmentation based on distinct premises, which rely on imagery only, labels only, or a combination of each. The proposed methods are able to consistently and effectively select the most important subset of samples (i.e., core-set) while filtering out non-representative and/or noisy samples.
Experiments were conducted using three high-resolution remote sensing datasets with very distinct properties: (i) IEEE GRSS Data Fusion Contest 2022 (DFC2022) dataset [25], consisting of very high-resolution visible spectrum images and Digital Elevation Model imagery, and (ii) Vaihingen and Potsdam datasets [26], both composed of high-resolution multispectral images and normalized Digital Surface Model data.
Experimental results have demonstrated the effectiveness and computational efficiency of the proposed techniques. Precisely, for the DFC2022, the proposed approaches outperformed the baseline trained on 100% of the data while using only 10% of the available samples. Similarly, for the Vaihingen and Potsdam datasets, the same superior performance was achieved using just 75% of the data. Overall, the core-set selection not only enhanced the performance of the deep learning models but also substantially reduced training time.
In summary, this work addressed a crucial gap in the literature and demonstrated the potential of core-set selection in advancing remote sensing image segmentation as well as data creation and labeling. The presented conclusions open opportunities towards a more efficient and effective exploitation of both existing and new datasets for a better understanding of the Earth’s surface, an essential characteristic for most applications.

### v2

@@ IV-B Implementation Details
We pre-process the aforementioned datasets by tiling them into non-overlapping $\times 256$ 256 patches that are used in all subsequent steps.
@@
All proposed methods [^1] The code is made publicly available at https://github.com/keillernogueira/data-centric-rs-classification/ are implemented using Pytorch. During training, we use the following hyper-parameters: 100 training epochs, AdamW [34] as optimizer, learning rate of 0.001, and batch size of 64 for the DFC2022 dataset and 32 for the other datasets.
@@
IV-C Baselines
For all datasets, we compare the proposed techniques with two traditional baseline models: (i) Random, which selects the core set uniformly at random, and (ii) CoreSet [12, 21, 22], which selects samples to optimally cover the embedding space. Precisely, this approach iteratively expands the core-set by adding the data point that is farthest from its nearest neighbor in the current (core) set. In this case, we used the Euclidean distance in the activations of the last spatial pooling layer of the ResNet-18 [28], similar to the approach used in the FA and FD methods.
IV-D Experimental Protocol
For the DFC2022 dataset, 90% of the data originally released for the data fusion contest is made available to be ranked by the proposed core-set algorithms, while the remaining 10% is used for validation.
For the Vaihingen and Potsdam datasets, we follow the standard protocol commonly exploited in the literature [35]. Specifically, for the Vaihingen dataset, 11 images originally released for the contest are made available for the proposed core-set techniques, and 5 images (with IDs 11, 15, 28, 30, 34) are employed for validation. For the Potsdam dataset, 18 images released for the contest are made available for the proposed techniques, whereas 6 images (with IDs 02_12, 03_12, 04_12, 05_12, 06_12, 07_12) are used for validation.
For all datasets, the validation is only employed to assess the training of the U-Net models, after the selection of the core-set. The final evaluation of the trained U-Net models uses the original test set of each dataset. The overall performance of each method is measured by the mean Intersection over Union (mIoU) across all segmentation classes and averaged over three different model training runs.
@@ V-A Quantitative Results
[…]
Table I shows the results for each method across the three datasets and different core-set sizes. Overall, the proposed approaches outperform the baselines for all datasets, even when trained using 100% of the available data. Moreover, in most cases, the proposed approaches outperform both baselines even when using substantially fewer training examples. For instance, in the Vaihingen dataset, the FA/CB Hybrid at 10% achieved a higher mIoU score than both baselines on the top-ranked 25% of the data. This shows the ability of the proposed techniques to select representative examples (core-set) for training.
@@
Finally, we also observe that each dataset has a different IoU convergence rate. DFC2022 has the most label noise (as evidenced by performance degradation of models in later splits), and the best-performing methods reached close to peak performance by utilizing only 10% of the data, noticeably outperforming training on 100% of the data. This rapid convergence suggests that for datasets with specific characteristics (e.g., redundancy, noise), a relatively small core-set can be as effective, or even more effective, than the full dataset. In contrast, on datasets with less label noise, such as Vaihingen and Potsdam, the performance continues to gradually improve as the number of training examples increases. However, even in this case, the introduced methods are able to outperform the baseline trained with 100% of the training data, demonstrating the importance of selecting a core set to deal with relevant issues such as noise, representativeness, and so on.
TABLE I: Results (mIoU) across different sized subsets of data (1%, 5%, 10%, 25%, 50%, 75%, and 100%) for DFC2022, Vaihingen, and Potsdam datasets. Underlined values indicate the results that outperformed the corresponding baselines per training percentage (statistically significant paired t-test at $\alpha=0.05$). Bold values represent the best results overall for the dataset.
@@
Category 1% 5% 10% 25% 50% 75% 100%
Random - 10.02 $\pm$ 0.37 11.02 $\pm$ 0.23 11.72 $\pm$ 0.12 12.38 $\pm$ 0.29 12.68 $\pm$ 0.11 12.02 $\pm$ 0.11 12.29 $\pm$ 0.12
CoreSet [12, 21, 22] - 10.58 $\pm$ 1.54 10.58 $\pm$ 1.54 11.08 $\pm$ 0.40 11.08 $\pm$ 0.90 11.50 $\pm$ 0.52 12.17 $\pm$ 0.34 12.10 $\pm$ 0.85
Label Complexity (LC) Label-only 10.54 $\pm$ 0.63 12.04 $\pm$ 0.20 12.68 $\pm$ 0.20 12.65 $\pm$ 0.03 12.46 $\pm$ 0.16 11.73 $\pm$ 0.76 -
Feature Diversity (FD) Image-only 10.71 $\pm$ 0.71 11.57 $\pm$ 0.44 11.77 $\pm$ 0.49 11.96 $\pm$ 0.05 12.17 $\pm$ 0.29 11.86 $\pm$ 0.27 -
LC/FD Hybrid Both 10.27 $\pm$ 0.26 12.43 $\pm$ 0.25 12.09 $\pm$ 0.50 12.54 $\pm$ 0.14 12.24 $\pm$ 0.34 12.28 $\pm$ 0.35 -
Feature Activation (FA) Image-only 9.14 $\pm$ 0.54 10.85 $\pm$ 0.21 10.86 $\pm$ 0.37 11.63 $\pm$ 0.18 12.32 $\pm$ 0.12 12.45 $\pm$ 0.26 -
Class Balance (CB) Label-only 9.12 $\pm$ 0.52 10.69 $\pm$ 0.90 11.01 $\pm$ 0.58 10.58 $\pm$ 0.19 11.95 $\pm$ 0.22 11.99 $\pm$ 0.19 -
FA/CB Hybrid Both 10.13 $\pm$ 0.10 10.14 $\pm$ 0.25 10.82 $\pm$ 0.22 11.98 $\pm$ 0.61 11.62 $\pm$ 0.07 12.10 $\pm$ 0.22 -
@@
Category 1% 5% 10% 25% 50% 75% 100%
Random - 36.20 $\pm$ 1.36 42.13 $\pm$ 0.52 50.72 $\pm$ 2.02 50.37 $\pm$ 2.97 55.82 $\pm$ 1.74 59.57 $\pm$ 0.64 59.95 $\pm$ 0.50
CoreSet [12, 21, 22] - 30.48 $\pm$ 1.49 32.65 $\pm$ 1.26 41.40 $\pm$ 0.54 45.68 $\pm$ 1.52 44.07 $\pm$ 3.85 49.17 $\pm$ 2.93 -
Label Complexity (LC) Label only 33.00 $\pm$ 4.03 45.80 $\pm$ 3.62 50.23 $\pm$ 4.33 53.75 $\pm$ 0.89 55.53 $\pm$ 0.41 58.70 $\pm$ 2.06 -
Feature Diversity (FD) Image only 33.67 $\pm$ 3.38 42.15 $\pm$ 0.57 44.87 $\pm$ 0.39 48.92 $\pm$ 3.60 56.07 $\pm$ 1.92 58.63 $\pm$ 0.50 -
LC/FD Hybrid Both 33.21 $\pm$ 3.50 40.45 $\pm$ 0.64 43.99 $\pm$ 0.42 51.15 $\pm$ 5.23 57.76 $\pm$ 0.46 58.97 $\pm$ 0.88 -
Feature Activation (FA) Image only 35.20 $\pm$ 1.22 43.61 $\pm$ 0.70 51.44 $\pm$ 0.69 53.44 $\pm$ 0.09 56.85 $\pm$ 0.60 60.58 $\pm$ 0.66 -
Class Balance (CB) Label only 40.92 $\pm$ 3.10 44.80 $\pm$ 1.23 51.03 $\pm$ 3.62 55.42 $\pm$ 0.83 58.36 $\pm$ 1.04 60.58 $\pm$ 0.30 -
FA/CB Hybrid Both 36.88 $\pm$ 1.39 49.06 $\pm$ 1.04 53.56 $\pm$ 0.25 56.92 $\pm$ 0.25 58.96 $\pm$ 0.67 60.43 $\pm$ 0.75 -
@@
Category 1% 5% 10% 25% 50% 75% 100%
Random - 60.89 $\pm$ 0.22 66.27 $\pm$ 0.88 68.21 $\pm$ 0.67 72.28 $\pm$ 0.66 74.56 $\pm$ 0.45 78.42 $\pm$ 0.33 78.00 $\pm$ 0.27
CoreSet [12, 21, 22] - 43.40 $\pm$ 7.30 42.69 $\pm$ 6.66 53.26 $\pm$ 6.27 68.85 $\pm$ 5.38 61.92 $\pm$ 1.28 64.98 $\pm$ 1.01 -
Label Complexity (LC) Label only 60.07 $\pm$ 2.33 68.68 $\pm$ 0.39 70.28 $\pm$ 0.45 75.12 $\pm$ 0.52 77.23 $\pm$ 0.79 79.15 $\pm$ 1.23 -
Feature Diversity (FD) Image only 59.78 $\pm$ 0.69 66.56 $\pm$ 0.35 69.37 $\pm$ 0.25 73.77 $\pm$ 0.54 75.42 $\pm$ 1.34 78.35 $\pm$ 1.21 -
LC/FD Hybrid Both 53.70 $\pm$ 6.05 68.25 $\pm$ 0.16 71.38 $\pm$ 0.34 74.87 $\pm$ 0.54 76.58 $\pm$ 0.56 78.72 $\pm$ 0.37 -
Feature Activation (FA) Image only 56.86 $\pm$ 1.47 65.75 $\pm$ 0.63 68.21 $\pm$ 0.58 72.89 $\pm$ 1.03 74.70 $\pm$ 0.52 79.58 $\pm$ 1.56 -
Class Balance (CB) Label only 49.89 $\pm$ 1.17 58.26 $\pm$ 0.39 61.94 $\pm$ 0.68 70.40 $\pm$ 0.30 75.29 $\pm$ 1.82 79.64 $\pm$ 0.48 -
FA/CB Hybrid Both 55.29 $\pm$ 1.49 61.72 $\pm$ 0.42 64.62 $\pm$ 0.40 71.06 $\pm$ 0.89 75.42 $\pm$ 0.78 78.49 $\pm$ 0.77 -
In addition to the gains in terms of performance, Table II reports the training time per epoch, in seconds, by a machine with an Intel Xeon E5-2695 v4 “Broadwell” CPU, 128GB of RAM memory, Nvidia P100 ”Pascal” GPU with 16GB of memory, under an 11.2 CUDA version, and Red Hat Enterprise Linux system release 8.2. As all models are trained for 100 epochs, this allows us to contextualize computational savings from the core-set methods. For example, with the DFC2022 dataset, the best-performing result (using 10% of the data) trained for 2 days and 2 hours less than the 100% training baseline while achieving a better test set performance. It is important to highlight that the core-set selection processing time is of the order of magnitude of one epoch with 100% of the data or less, making it negligible compared to the full 100-epoch training. Additionally, faster training time allows for more epochs and/or more extensive hyperparameter tuning, further enhancing model optimization.
@@ V-B Qualitative Results
To facilitate the analysis and comparison of the proposed methods’ outputs, we include visualizations of the average generated rankings, as can be seen in Figure 4. To generate these visualizations, the rankings produced by the different introduced methods are first averaged by patch position and then sorted, making the highest-ranking patches across the methods appear at the top. Additionally, the standard deviation is calculated to capture the variability of the assigned ranks and provide insight into the approaches’ consistency. In addition to these visualizations, to allow for a better analysis, we also report the Kendall Tau correlations between each pair of methods in Figure 5 and provide examples of the most and least frequently selected instances across all introduced methods in Figure 6.
@@
Figure 4: Visualizations of the proposed methods’ rankings. The line represents the average rank position for each patch across all proposed approaches, while the shaded area represents the standard deviation.
@@
Figure 5: Correlation of methods according to Kendall Tau coefficient. A high correlation value means that the methods produce similar rankings.
For all datasets, it is possible to observe that the approaches exhibit notable consistency, frequently assigning more importance to a specific (core) set of high-complexity examples that are consistently selected across the proposed methods. Similarly, such approaches also tend to agree on the least important examples, assigning lower scores to low-complexity patches, indicating that the explored datasets contain a subset of non-representative or noisy instances that either contribute minimally to the overall performance or, in some cases, may even degrade it. Furthermore, this level of agreement between the proposed techniques can be further observed in the correlation plots, wherein several methods show substantial correlation (particularly for the DFC2022 and Potsdam datasets), suggesting that they can capture underlying dataset patterns (such as the core sets). Overall, the ability to select the core set, along with the identification of less valuable examples, highlights the robustness and efficiency of the proposed techniques in distinguishing between high- and low-quality data, thereby resulting in better performance and training time. This is also qualitatively demonstrated in Figure 6: examples that are consistently highly ranked by the proposed methods show clear imagery with artifacts that are of a certain visual and semantic complexity (as illustrated by the corresponding label maps). A few of these images allow learning the appearance and spatial relation of several classes at once. On the other hand, images that are consistently rejected show very homogeneous scenes (such as large water bodies or parking lots) with neither much visual variation nor complex semantic content.
@@
Figure 6: Examples of the highest-ranked (first three columns) and lowest-ranked instances (last three columns) considering the average ranking of all proposed methods.
@@ VI Conclusions
In this paper, we introduce and benchmark six basic core-set selection approaches for remote sensing image segmentation based on distinct premises - which rely on imagery only, labels only, or a combination of each - thereby establishing a general and comprehensive baseline for future works. The proposed methods are able to consistently and effectively select the most important subset of examples (i.e., core-set) while filtering out non-representative and/or noisy instances.
Experiments are conducted using three high-resolution remote sensing datasets with very distinct properties: (i) IEEE GRSS Data Fusion Contest 2022 (DFC2022) dataset [31], consisting of very high-resolution visible spectrum images and Digital Elevation Model imagery, and (ii) Vaihingen and Potsdam datasets [32], both composed of high-resolution multispectral images and normalized Digital Surface Model data.
Experimental results demonstrate the effectiveness and computational efficiency of the proposed techniques, which consistently outperform the baselines. Notably, on the DFC2022 dataset, the proposed approaches outperform the baseline trained on 100% of the data while using only 10% of the available examples. Similarly, on the Vaihingen and Potsdam datasets, the same superior performance is achieved using just 75% of the data. Overall, the core-set selection not only enhances the performance of the deep learning models but also substantially reduces training time.
In summary, this work addresses a crucial gap in the literature and demonstrates the potential of core-set selection in advancing remote sensing image segmentation as well as data creation and labeling. The presented conclusions open opportunities towards: (i) the integration of core-set selection with other advanced techniques, such as self-supervised learning and foundation models, and (ii) a more efficient and effective exploitation of both existing and new datasets for a better understanding of the Earth’s surface, an essential characteristic for most applications.